OpenStudy (anonymous):

i need to know how to solve and the answer for E over n=1 40 (3/5)^n-1 and i need it right away I gave like two other questions like this and I need to pass this course

OpenStudy (anonymous):

this is the problem and I don't know how to do it

OpenStudy (anonymous):

Karp

OpenStudy (anonymous):

From Mathematica v9:\[\sum _{n=1}^{\infty } \frac{3}{8} \left(\frac{3}{4}\right)^{n-1}=\frac{3}{2} \]